A 10g riffle bullet acquires a speed of 400m/s in traversing a barrel 50cm long. Find the average

acceleration and the accelerating force.


1
Expert's answer
2022-03-29T16:46:24-0400

The average acceleration

"a=\frac{v_f^2-v_i^2}{2d}=\frac{0^2-400^2}{2*0.5}=-1.6*10^5\:\rm m/s^2"

The magnitude of the force

"F=ma\\
=0.01*1.6*10^5=1600\:\rm N"